OpenStudy (anonymous):

2. 2^3x+2=64 1. 8/3 2. 4/3 3. 2 4. 3/4 3.log(3x+25)=2 1.25 2.75 3.2/3 4.100 4.16^2x=124 1.0.869 2.1.150 3.1.739 4.3.477

OpenStudy (campbell_st):

can I just check the question is \[2^{3x +2} = 64\]

OpenStudy (aravindg):

is it ?

OpenStudy (anonymous):

Yes. if theres a ^ that part goes up

OpenStudy (campbell_st):

ok.. for the 1st question 64 can be written as 2^6 so you are being asked to solve \[2^{3x + 2} = 2^6\] all you need to do is equate the powers 3x + 2 = 6 now just solve for x

OpenStudy (anonymous):

x=1.3

OpenStudy (campbell_st):

I'd write it as x = 4/3 since your answer choices are fractions.